Determine the equation of the straight line which passes through the intersection of the lines given by 3x-4y+1=0 and 5x+y=1 and has equal intercepts of the same sign along the axes
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
on solving these lines
x = 3/23
y = 8/23
 
now the line make equal intercept hence m = tan 135 = -1
 
hence equation of line 
y – 8/23 = – 1 (x- 3/23)